OpenStudy (anonymous):

evaluate the function at x=-4, x=-5. Write undefined if appropriate. x-6 y= -------- x^2+7x+10

OpenStudy (anonymous):

(-5)^2+7(-5)+10 = 25-35+10 =35-35=0 Devision by zero is undefined

OpenStudy (anonymous):

Division

jhonyy9 (jhonyy9):

- so yes for x=-5 the denominator will be equal zero and we know that than the fraction is undefined sure

jhonyy9 (jhonyy9):

ok ?